OpenStudy (anonymous):

use the change of base formula to evaluate log3 58 to a logarithm in base 4. round to the nearest thousandth. Explain STEP BY STEP HOW YOU GOT YOUR ANSWER

OpenStudy (anonymous):

how the hell are you going to find the decimal approximation if you use log base 4? you do not have log base 4 on your calculator, only base 10 and base \(e\)

OpenStudy (anonymous):

if you want a number out of this, you need \[\log_3(58)=\frac{\ln(58)}{\ln(3)}\]

OpenStudy (anonymous):

convert log3 58 to a logarithm in base 4 and round to the nearest thousandth

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

you can convert it to log base 4 by writing \[\frac{\log_4(58)}{\log_4(3)}\] but that doesn't help your cause as far as finding an actual number for an answer

OpenStudy (anonymous):

@satellite73 i have the answer i just need to know the steps to getting it

OpenStudy (anonymous):

@satellite73